From GlueXWiki
Jump to: navigation, search

October 1, 2015 Drift Chamber meeting


  1. Instructions for Bluejeans meeting connection
  2. Meeting ID: 290664653
  3. To join via a Web Browser, go to the page [1] https://bluejeans.com/290664653.


  1. CDC update (Mike, Simon)
  2. FDC update (Luke, Lubomir)
  3. Electronics (Fernando, Chris, Nick)
  4. Engineering (Dave)
  5. Other


Participants: Curtis, Naomi, Mike (CMU), Sean (NU), Simon, Luke, Eugene, Marc, Beni, Dave, Chris, Sergey, Cody and Lubomir (JLab).

CDC update

- Simon presented a report (linked) about parametrization of the time-to-distance for different offsets (delta) of the wire from the center of the tube. The time-to-distance is fitted with a function which parameters depend on delta linearly. This works for big times, for short he uses tables and a hybrid model in between. Maximum drift time (for a few tubes only) can be up to 2 microseconds.

- Mike presented another report on studying the offsets with cosmics, offering better phi coverage than with beam tracks originating from the target. The results show mostly downward deformations of the tubes (sagging) at most 2mm with a few exceptions up to 3mm. The deformations are grouped in layers mostly for the stereo ones.

FDC update

- Luke showed plots for the avalanche positions with 55Fe source, this time using the max amplitude instead of the integral. The results are very similar to the previous ones.


- Cody on the f125 firmware: introduced a hold-off if the buffer is full to avoid crushing in case several events come within short time. Will test it with Sergey. Now the max window size is 300 (?) samples, still can have a different version for special cases like the one for the TRD studies.


-Lubomir tested the FDC MFCs in the Hall. The MFC for the second package showed the similar problem as the first one (see last meeting) - it stops regulating the flow (set at 200ccpm) when the input pressure goes below ~8 psi (the first stopped regulating below 10 psi). Eugene suggested decreasing the variations of the pressure in the mixing tank from 7-12 to say 10.8 -12 psi. The last two MFC are OK. Beni wanted to have their serial numbers and look back - the MFCs were send for repair/recalibration at different times.


- No meeting next week due to the collaboration meeting.